solve each of the following quadratic inequalities. express your answers in interval notation.
a) 2y^2+5y-63<0

b) 4y^2+10y-14<0
c) x^2+x-30<0

Answer by venugopalramana(3286)

SEE THE FOLLOWING EXAMPLE AND TRY.IF STILL IN DIFFICULTY PLEASE COME BACK
-------------------------------
PLEASE HELP ME WITH THIS PROBLEM;
For f(x)=1 over x62-2x-8 find the intervals where f(x)<0.
HOPE YOU MEAN
Y = F(X)=1/ X^2-2X-8.....ASSUMING SO
Y= 1/X^2-4X+2X-8=1/[X(X-4)+2(X-4)]=1/[(X+2)(X-4)]
THE ROOTS ARE X=-2 AND X=4
FOR F(X) TO BE <0,X SHOULD LIE BETWEEN THE 2 ROOTS ..THAT IS ..BETWEEN -2 & 4 .
-2 -------------------------------------------------------
ANSWERS ARE
a) 2y^2+5y-63<0 ...(2Y-9)(Y+7)<0....Y LIES BETWEEN -7 AND 9/2...(-7,9/2)
b) 4y^2+10y-14<0 ...2(2Y+7)(Y-1)<0...Y LIES BETWEEN -7/2 AND 1 ...(-7/2,1)
c) x^2+x-30<0....(X+6)(X-5)<0....X LIES BETWEEN -6 AND 5......(-6,5)
You may edit the question. M